find the equation of a tangent line to ellipse 9x^2+25y^2=225 at x=3

Grade:12th pass

1 Answers

Rinkoo Gupta
9 years ago
put x=3 in eq. of the ellipse
9(9)+25y2=225
81+25y2=225
25y2=144
y2=144/25
y=12/5
so the touching point is (3,12/5)
eq. of the ellipse can be written as x2/25 +y2/9 =1
noq eq of the tangent to the ellipse at point (3,12/5) is
3x/25 +12y/45=1
27x+60y=225 Ans.
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty